# Security and Privacy Challenges of Large Language Models: A Survey

### v1

@@ 5.3. PII Leakage Attack
PII, refers to data that, either alone or in combination with other information, can uniquely identify an individual (phishing, ). PII encompasses direct identifiers like passport details and quasi-identifiers such as race and date of birth. Sensitive PII includes information like name, phone number, address, social security number (SSN), financial, and medical records, while non-sensitive PII, is readily available in public sources, such as zip code, race, and gender. Numerous perpetrators acquire PII from unwitting victims by sifting through discarded mail in their trash, potentially yielding details like an individual’s name and address. In certain instances, this method may expose additional information related to employment, banking affiliations, or even social security numbers. Phishing and social engineering attacks (gupta2016literature, ) leverage deceitful websites or emails, employing tactics designed to deceive individuals into disclosing critical details such as names, bank account numbers, passwords, or SSNs. Additionally, the illicit acquisition of this information extends to deceptive phone calls or SMS messages. In LLMs, PII leakage has been a fundamental problem. In March 2023, it was reported that ChatGPT leaked users’ conversation history as well as information related to payment due to a bug in the system (kshetri2023cybercrime, ). Evidence has been found on leaking information through sentence-level MIA (shokri2017membership, ) and reconstruction attacks (zheng2023input, ) on private training data. One of the first studies of PII leakage was proposed by Inan et al. named TAB attack (inan2021training, ). Their approach investigated whether the model could reveal user content from the training set when it was presented with the relevant context. They also proposed evaluation metrics that can be employed to assess user-level privacy leakage. After that, Lukas et al. empirically demonstrated that their attack method against GPT-2 models can extract up to $10\times$ more PII sequences than TAB attack. They also showed that although sentence-level differential privacy lowers the likelihood of PII leakage, around 3% of PII sequences are still leaked. PII reconstruction and record-level membership inference were shown to have a subtle relationship (lukas2023analyzing, ). Zanella et al. (zanella2020analyzing, ) explored the impact of updates on language models by analyzing snapshots before and after an update, revealing insights into changes in training data. Two metrics were introduced by them, differential score, and differential rank, to assess data leakage in natural language models, which includes a privacy analysis of language models trained on overlapping data, demonstrating that adversaries can extract specific content without knowledge of training data or model architecture. ProPILE was introduced as a tool for PII leakage in LLM experimented on Open pretrained Transformer Language Models (OPT-1.3B model (zhang2022opt, )). It will ask for a specific PII, e.g., contact no. or SSN to the designed LLM prompt by providing associated information. Then the LLM prompt will provide the asked information based on the likelihood of that given information formulated by linkability of the given information and structure of the asked information (kim2023propile, ). Researchers addressed the PII learning tasks of LLMs and showed that PII that was forgotten might be retrieved by fine-tuning using a few training instances (chen2023janus, ). Considering some primary factors of privacy leakage in PLMs, a universal framework named KART has been introduced specifically for the biomedical domain (nakamura2020kart, ).
Memorization is another aspect of PII attacks. Carlini et al. demonstrated that LLMs memorize and leak individual training examples (carlini2021extracting, ). Additionally, it shows how a malicious party may query the language model in order to execute a training data extraction attack and get specific training samples (GPT-2). It showed that LLMs memorize and leak individual training examples. A straightforward approach was proposed to use only black-box query access, where verbatim sequences (as exactly they appeared in the training set) were extracted from a language model’s training set (carlini2021extracting, ). It can be directly applied to any language model trained on non-public and sophisticated data. The GPT-2 model released by OpenAI has been a representative language model used in the experiments. Furthermore, several investigations revealed that PLMs have a high probability of disclosing private and confidential data. In particular, if it asks PLMs for email addresses along with email address contexts or asks for prompts that include the owner’s name. According to the studies, PLMs retain personal data, which means that the data may be retrieved using a certain prefix, such as training data tokens. PLMs link the owner of the personal information to it, thus attackers may query the data using the owner’s identity (huang2022large, ).
@@ 6. Defense Mechanisms
As LLMs become integral components in applications ranging from NLP to multimodal systems, the vulnerabilities associated with their usage pose serious concerns. Protecting LLMs from security and privacy attacks is imperative to preserve the reliability and integrity of this complex AI system (defNdd, ). We argue that robust defense strategies should be developed to safeguard LLMs from both security and privacy perspectives. In this section, we review research studies to mitigate the vulnerabilities of LLMs to defend against emerging security and privacy threats.
@@ 6.1. Defense Against Security Attacks on LLMs
Defense Against Prompt Injection. Limited studies explored the defense strategies to defend the prompt injection attacks in LLMs. A prevention-detection-based defense technique has been reported to systematically present existing defense mechanisms against prompt injection attacks (liu2023prompt, ). Prevention-based defenses, as outlined in (Prmt111, ) and (jain2023baseline, ), are designed to thwart the successful execution of tasks injected into an LLM-integrated application. These preventive measures involve preprocessing the data prompt to eliminate the injected task’s instruction/data, and/or redesigning the instruction prompt itself. To thwart the adversarial prompts there are several techniques, e.g., paraphrasing (jain2023baseline, ), re-tokenization (jain2023baseline, ), data prompt isolation, and instructional prevention. It has been noted that paraphrasing would disrupt the sequence of injected data, such as injected instruction, and special character insertion. The efficacy of prompt injection attacks would be diminished by this disruption. Re-tokenization aims to break the sequence of injected instructions, task-ignoring text, special characters, and fake responses within a compromised data prompt. This process preserves frequently occurring words while breaking down infrequent ones into multiple tokens. Consequently, the re-tokenized output comprises more tokens than a typical representation. This re-tokenized data prompt and the instruction prompt are used by the LLM-Integrated application to query the LLM and generate a response. Defenses based on detection are focused on determining the integrity of a given data prompt (jain2023baseline, ), (prmttt, ), (socher2013recursive, ), (wang2022perplexity, ). Notably, the proactive detection method (prmttt, ) has proven effective in identifying instances of prompt injection attacks. Defenses based on detection can further be classified into two categories: response-based detection, and prompt-based detection. A response-based detection method examines the response of LLMs, while a prompt-based detection approach examines a provided data prompt. Perplexity-based detection is a kind of prompt-based detection. The basic idea is that adding information or instructions to a data prompt degrades its quality and leads to increased perplexity. Consequently, a data prompt is considered compromised if its perplexity exceeds a specified threshold (gonen2022demystifying, ). Since an LLM-integrated application is tailored for a specific task, granting it prior knowledge of the anticipated response, detecting a compromised data prompt is feasible when the generated response deviates from a proper answer for the desired task (prmttt, ). For example, if the desired task is spam detection and the response does not align with “spam” or “non-spam”, they imply a compromise. Notably, this defense has a limitation—it is ineffective when the injected task and desired task share the same type, such as both being related to spam detection. Effective protection strategies against P2SQL injection attacks are available and may be incorporated into the Langchain framework as extensions (pedro2023prompt, ). For example, database permission hardening, since P2SQL injection attacks can manipulate chatbots by arbitrarily executing different queries (pedro2023prompt, ), including deleting data, utilizing database roles and permissions to limit the execution of undesired SQL statements when accessing tables with sensitive information can be a viable technique to defend such attacks. It can mitigate arbitrary access by rewriting the SQL query output by LLM into a semantically equivalent one that exclusively operates on the information the user is authorized to access (pedro2023prompt, ). Auxiliary LLM Guard is another way to mitigate the P2SQL attacks. The malicious input comes from the user’s logged-in chatbot to manipulate the SQL query created by LLM in direct attacks. Conversely, indirect attacks involve malicious input residing in the database, enabling interference with LLM-generated SQL queries and potentially undermining the effectiveness of these defenses. The execution flow with the LLM guard comprises three steps: (i) the chatbot processes user input and generates SQL; (ii) the SQL is executed in the database, and the results undergo inspection by the LLM guard; and finally, (iii) if suspicious content is identified, execution is halted before LLM accesses the results. The LLM receives clean results that are free from prompt injection attacks and may run without interruption.
@@
Defense Against Jailbreaking Attacks. Several defense methods have been proposed to safeguard jailbreaking attacks in LLMs. As a built-in safety mechanism, preprocessing-based techniques, detecting and blocking the inputs or outputs, and semantic content filtering have been employed to prevent generating undesired or inappropriate contents from LLMs, which could effectively mitigate potential harm (markov2023holistic, ). Kupmar et al. (kumar2023certifying, ) proposed an approach to apply a safety filter on the sub-strings of input prompts, which provides certifiable robustness guarantees. The drawback of this approach lies in the method’s complexity, which increases proportionally with input prompt length. Wu et al. (wu2023defending, ) propose a system-mode self-reminder to defend against jailbreaking attacks under the pretending or role-playing scenarios, which can drastically reduce the jailbreaking success rate from 67.21% to 19.34%. It is a technique to assist ChatGPT in remembering or focusing on particular actions, ideas, or behaviors when it is asked to generate inappropriate content (wu2023defending, ). One potential simple defense strategy is to identify the presence of “red-flagged” keywords (wei2023jailbroken, ) which strictly violates the usage policies of the LLM vendors, e.g., OpenAI (deng2023jailbreaker, ). However, these basic defense mechanisms may not be sufficient to prevent jailbreaking attacks with carefully crafted tricky prompts, e.g., privilege escalation. Furthermore, preventing these attacks still poses a significant challenge because the effective defenses may impair model utility (li2023sok, ). By far, SmoothLLM is an effective defense strategy against existing jailbreaking attacks proposed by Zou et al. (zou2023universal, ). The fundamental concept of SmoothLLM is partially inspired by the randomized smoothing within the adversarial robustness community (cohen2019certified, ), involving a two-step process. First, copies of a specified input prompt are duplicated and perturbed. Subsequently, the outputs produced for each perturbed copy are aggregated. The desiderata encompass four key properties: attack mitigation (reduces attack success rate 100 times and 50 times for Llama2 and Vicuna respectively), non-conservatism, efficiency (in terms of computational resources), and compatibility (different LLMs). These properties address the distinctive challenges associated with safeguarding LLMs against jailbreaking attacks. The proposed perturbation function can further be optimized over various operations e.g., insertion and swaps to make stronger defenses. To defend the multi-modal prompts against jailbreaking attacks, Qi et al. recently proposed DiffPure (nie2022diffusion, ), a diffusion model-based countermeasure against the visual jailbreaking examples.
@@
Defense Against Backdoor Attack. Most of the existing research, such as the removal of backdoors by fine-tuning (sha2022fine, ), model pruning (liu2018fine, ), and detecting backdoors by inspecting activations (chen2018detecting, ) are based on backdoor defenses in the white-box setting. Fine-mixing is a mitigation approach designed to prevent backdoors in fine-tuned LMs. It utilizes pretrained weights through two complementary techniques: (i) a two-step fine-tuning procedure that first combines backdoored weights that have been optimized using pretrained weights on poisoned data, and then refines the combined weights on a small collection of clean data; (ii) an Embedding Purification (E-PUR) method, addressing potential backdoors in word embeddings (zhang2022fine, ). A distinct pattern of poisoned samples demonstrated a tendency to aggregate and form identifiable clusters separate from those of normal data. Building upon this observation, a defense technique named CUBE has been proposed (cui2022unified, ). It utilized a density clustering algorithm called HDBSCAN to accurately discern clusters within datasets, distinguishing between those containing poisoned samples and those with clean data. By leveraging the capabilities of HDBSCAN (mcinnes2017hdbscan, ), CUBE aims to provide an effective means of differentiating clusters associated with both normal and poisoned data. Strategies to defend the backdoor attacks in the black box setting are still lacking (kandpal2023backdoor, ). Perturbation-based and perplexity-based defense methods are also adopted in the literature (mozes2023use, ). Several existing works have been designed where the users fine-tune a model on their custom clean data, e.g., RAP which leverages word-based robustness-aware perturbation to identify poisoned samples (yang2021rap, ) and ONION which eliminates trigger words via empirical analysis of sentence perplexities (qi2020onion, ). Masking-Differential Prompting (MDP) serves as an efficient, lightweight, and adaptable defense method against backdoor attacks in prompt-based language models (PLMs), particularly in few-shot learning scenarios (xi2023defending, ). MDP exploits the observation that poisoned samples exhibit increased sensitivity to random masking compared to clean samples. When the trigger is (partially) masked, the language modeling probability of a poisoned sample tends to exhibit significant variations. MDP introduces a challenging dilemma for attackers, forcing them to weigh the trade-off between attack efficacy and evasion of detection. However, MDP falls short on several other PLMs (e.g., GPT-3 (brown2020language, )) and NLP tasks (e.g., paraphrases and sentence similarity (gao2020making, )). Also, it was not evaluated under large few-shot data in practice, therefore, it is still unclear how it would react to the large few-shot data. Again, MDP was proven to be effective for the earlier backdoor attacks, however, it might not safeguard some advanced attacks such as BadPrompt (cai2022badprompt, ) and BToP (xu2022exploring, ).
@@
Defense Against Data Poisoning Attack. Few solutions exist to defend against poisoning attacks in LLMs. In general, techniques such as data validation, filtering, cleaning, and anomaly detection have been used to protect ML models from poisoning attacks (mlpoision, ). A detection and filtering approach was designed to identify and filter poisoned data which is collected for performing supervised learning (baracaldo2017mitigating, ). Empirical assessments have demonstrated that limiting the number of training epochs is a straightforward method for LMs to reduce the impact of data poisoning such as RoBERTa (wallace2020concealed, ). Identifying poison examples using perplexity can be another technique for small GPT-2 model (radford2019language, ) for sentiment analysis tasks. However, it may not identify poisons effectively, specifically, after inspecting the training data, less than half of the poisoned examples can be identified. Identifying these poisoned examples by BERT embedding distance is another method for defending this attack (wallace2020concealed, ). Filtering poisoned samples during training is also used in LMs to defend against data poisoning attacks. The poisoned data points are often outliers in the training data distribution. Compared to normal benign training data, poisoned data require more time to enable the model to learn their features. To defend against trojaning attacks in LMs, like, TROJANLM (zhang2021trojaning, ), one possible defense technique is to adopt existing techniques from other domains such as images, e.g., detecting trigger-embedded inputs at inference time (chen2018detecting, ), (chou2018sentinet, ), and finding suspicious LMs and retrieving triggers during the model evaluation phase (wang2019neural, ), (chen2019deepinspect, ). Dataset cleaning techniques, such as removing near-duplicate poisoning samples, known triggers and payload, and removing anomalies can be potentially applied to defend against TROJANPUZZLE attack (aghakhani2023trojanpuzzle, ).
@@
The aforementioned defense strategies are mostly designed for LMs, but some of them can be potentially applied to LLMs as well, however, it still lacks in-depth research studies to develop efficient defense techniques to protect LLMs from data poisoning attacks. Moreover, empirical reports have shown that LLMs are becoming more vulnerable to data poisoning attacks, where defenses based on filtering data or lowering model capacity only offer minimal protection at the cost of reduced test accuracy (wan2023poisoning, ). Therefore, it requires effective defense methods that can trade-off between the model utility and the capability of protecting LLMs from data poisoning attacks.

### v2

@@ 5.3. PII Leakage Attack
PII, refers to data that, either alone or in combination with other information, can uniquely identify an individual (FRANKENFIELD, 2023). PII encompasses direct identifiers like passport details and quasi-identifiers such as race and date of birth. Sensitive PII includes information like name, phone number, address, social security number (SSN), financial, and medical records, while non-sensitive PII, is readily available in public sources, such as zip code, race, and gender. Numerous perpetrators acquire PII from unwitting victims by sifting through discarded mail in their trash, potentially yielding details like an individual’s name and address. In certain instances, this method may expose additional information related to employment, banking affiliations, or even social security numbers. Phishing and social engineering attacks (Gupta et al., 2016) leverage deceitful websites or emails, employing tactics designed to deceive individuals into disclosing critical details such as names, bank account numbers, passwords, or SSNs. Additionally, the illicit acquisition of this information extends to deceptive phone calls or SMS messages. In LLMs, PII leakage has been a fundamental problem. In March 2023, it was reported that ChatGPT leaked users’ conversation history as well as information related to payment due to a bug in the system (Kshetri, 2023). Evidence has been found on leaking information through sentence-level MIA (Shokri et al., 2017) and reconstruction attacks on private training data (Zheng, 2023). One of the first studies of PII leakage was proposed by Inan et al. named TAB attack (Inan et al., 2021b). Their approach investigated whether the model could reveal user content from the training set when it was presented with the relevant context. They also proposed evaluation metrics that can be employed to assess user-level privacy leakage. After that, Lukas et al. empirically demonstrated that their attack method against GPT-2 models can extract up to 10×10$\times$10 × more PII sequences than TAB attack. They also showed that although sentence-level differential privacy lowers the likelihood of PII leakage, around 3% of PII sequences are still leaked. PII reconstruction and record-level membership inference were shown to have a subtle relationship (Lukas et al., 2023). Zanella et al. (Zanella-Béguelin et al., 2020) explored the impact of updates on LMs by analyzing snapshots before and after an update, revealing insights into changes in training data. Two metrics were introduced by them, differential score, and differential rank, to assess data leakage in natural language models, which includes a privacy analysis of LMs trained on overlapping data, demonstrating that adversaries can extract specific content without knowledge of training data or model architecture. ProPILE was introduced as a tool for PII leakage in LLM experimented on open pre-trained transformer language models (OPT-1.3B model (Zhang et al., 2022b)). It will ask for a specific PII, e.g., personal contact number or SSN to the designed LLM prompt by providing associated information. Then the LLM prompt will provide the asked information based on the likelihood of that given information formulated by linkability of the given information and structure of the asked information (Kim et al., 2024). Researchers addressed the PII learning tasks of LLMs and showed that forgotten PII might be retrieved by fine-tuning using a few training instances (Chen et al., 2023). Considering some primary factors of privacy leakage in PLMs, a universal framework named KART has been introduced specifically for the biomedical domain (Nakamura et al., 2020). Nakka et al. proposed PII leakage attack, termed PII-compass (Nakka et al., 2024) by adding hand-crafted template to an attack prefix. These attack prefixes are associated with a different data subject from the adversary set, meaning the data (attack prefix) used to create the new prompt originates from a different data subject whose PII is intended for extraction. Memorization is another aspect of PII attacks. Carlini et al. demonstrated that LLMs memorize and leak individual training examples (Carlini et al., 2021). Additionally, it shows how a malicious party may query the LM in order to execute a training data extraction attack and get specific training samples (GPT-2). It showed that LLMs memorize and leak individual training examples. A straightforward approach was proposed to use only black-box query access, where verbatim sequences (as exactly they appeared in the training set) were extracted from a language model’s training set (Carlini et al., 2021). It can be directly applied to any LM trained on non-public and sophisticated data. The GPT-2 model released by OpenAI has been a representative LM used in the experiments. Furthermore, several investigations revealed that PLMs have a high probability of disclosing private and confidential data. In particular, when it asks PLMs for email addresses along with email address contexts or asks for prompts that include the owner’s name. According to the studies, PLMs retain personal data, which means that the data may be retrieved using a certain prefix, such as training data tokens. PLMs link the owner of the personal information to it, thus attackers may query the data using the owner’s identity (Huang et al., 2022).
Certain shortcomings exist in the PII leakage attacks prevalent in the literature. For instance, Huang et al. primarily focused on extracting email addresses as representatives of personal information in their proposed memorization attack in PLM (Huang et al., 2022). On the contrary, PII-compass (Nakka et al., 2024) is limited to only extracting phone numbers. Complete extraction of a full PII set still requires significant exploration and proper methods. While some methods have been evaluated only on small Transformer-based language models (Inan et al., 2021b), others have been tested on earlier LLMs such as GPT-2 (Carlini et al., 2021). The performance of these methods (e.g., (Inan et al., 2021b; Huang et al., 2022)) on the latest LLMs, such as Llama-3 and GPT-4, still requires in-depth evaluation. Moreover, some of these methods may not perform well in the presence of defense mechanisms such as differential privacy (DP) (Abadi et al., 2016). Therefore, further research is needed to develop attack methods that address the aforementioned issues in PII leakage attacks against LLMs.
@@ 6. Defense Mechanisms
As LLMs become integral components in applications ranging from NLP to multi-modal systems, the vulnerabilities associated with their usage pose serious concerns. Protecting LLMs from security and privacy attacks is imperative to preserve the reliability and integrity of this complex system (OpenAI, 2023a). We argue that robust defense strategies should be developed to safeguard LLMs from security and privacy perspectives. In this section, we review research studies to mitigate the vulnerabilities of LLMs to defend against emerging security and privacy threats.
@@ 6.1. Defense Against Security Attacks on LLMs
Defense Against Prompt Injection. Limited studies explored the defense strategies to defend the prompt injection attacks in LLMs. A prevention-detection-based defense technique has been reported to systematically present existing defense mechanisms against prompt injection attacks (Liu et al., 2023a). Prevention-based defenses, as outlined in (Prompting, 2023b) and (Jain et al., 2023), are designed to thwart the successful execution of tasks injected into an LLM-integrated application. These preventive measures involve pre-processing the data prompt to eliminate the injected task’s instruction/data, and/or redesigning the instruction prompt itself. To thwart the adversarial prompts there are several techniques, e.g., paraphrasing (Jain et al., 2023), re-tokenization (Jain et al., 2023), data prompt isolation, and instructional prevention (Schulhoff, 2024a, b, c). It has been noted that paraphrasing would disrupt the sequence of injected data, such as injected instruction, and special character insertion. The efficacy of prompt injection attacks would be diminished by this disruption. Re-tokenization aims to break the sequence of injected instructions, task-ignoring text, special characters, and fake responses within a compromised data prompt. This process preserves frequently occurring words while breaking down infrequent ones into multiple tokens. Consequently, the re-tokenized output comprises more tokens than a typical representation. This re-tokenized data prompt and the instruction prompt are used by the LLM-Integrated application to query the LLM and generate a response. Defenses based on detection are focused on determining the integrity of a given data prompt (Jain et al., 2023), (Selvi., 2023), (Socher et al., 2013), (Wang et al., 2022a). Notably, the proactive detection method (Selvi., 2023) has proven effective in identifying instances of prompt injection attacks. Defenses based on detection can further be classified into two categories: response-based detection, and prompt-based detection. A response-based detection method examines the response of LLMs, while a prompt-based detection approach examines a provided data prompt. Perplexity-based detection is a kind of prompt-based detection. The basic idea is that adding information or instructions to a data prompt degrades its quality and leads to increased perplexity. Consequently, a data prompt is considered compromised if its perplexity exceeds a specified threshold (Gonen et al., 2022). Since an LLM-integrated application is tailored for a specific task, granting it prior knowledge of the anticipated response, detecting a compromised data prompt is feasible when the generated response deviates from a proper answer for the desired task (Selvi., 2023). For example, if the desired task is spam detection and the response does not align with “spam” or “non-spam”, they imply a compromise. Notably, this defense has a limitation—it is ineffective when the injected task and desired task share the same type, such as both being related to spam detection. Effective protection strategies against P2SQL injection attacks are available and may be incorporated into the LangChain framework as extensions (Pedro et al., 2023). For example, database permission hardening, since P2SQL injection attacks can manipulate chatbots by arbitrarily executing different queries (Pedro et al., 2023), including deleting data, utilizing database roles and permissions to limit the execution of undesired SQL statements when accessing tables with sensitive information can be a viable technique to defend such attacks. It can mitigate arbitrary access by rewriting the SQL query output by LLM into a semantically equivalent one that exclusively operates on the information the user is authorized to access (Pedro et al., 2023). Auxiliary LLM Guard is another way to mitigate the P2SQL attacks. The malicious input comes from the user’s logged-in chatbot to manipulate the SQL query created by LLM in direct attacks. Conversely, indirect attacks involve malicious input residing in the database, enabling interference with LLM-generated SQL queries and potentially undermining the effectiveness of these defenses. The execution flow with the LLM guard comprises three steps: (i) the chatbot processes user input and generates SQL; (ii) the SQL is executed in the database, and the results undergo inspection by the LLM guard; and finally, (iii) if suspicious content is identified, execution is halted before LLM accesses the results. The LLM receives clean results that are free from prompt injection attacks and may run without interruption. Liu et al. performed an empirical study to comprehensively formalize and benchmark several prompt injection attacks and defenses (Liu et al., 2024a). In their study, they found the prevalent prevention-based and detection-based defense techniques are insufficient to mitigate the risks of advanced optimization-based attacks, e.g., JudgeDeceiver (Shi et al., 2024).
@@
Defense Against Jailbreaking Attacks. Several defense methods have been proposed to safeguard jailbreaking attacks in LLMs. As a built-in safety mechanism, pre-processing-based techniques, detecting and blocking the inputs or outputs, and semantic content filtering have been employed to prevent generating undesired or inappropriate contents from LLMs, which could effectively mitigate potential harm (Markov et al., 2023). Kupmar et al. (Kumar et al., 2023) proposed an approach to apply a safety filter on the sub-strings of input prompts, which provides certifiable robustness guarantees. The drawback of this approach lies in the method’s complexity, which increases proportionally with input prompt length. Wu et al. (Wu et al., 2023c) propose a system-mode self-reminder to defend against jailbreaking attacks under the pretending or role-playing scenarios, which can drastically reduce the jailbreaking success rate from 67.21% to 19.34%. It is a technique to assist ChatGPT in remembering or focusing on particular actions, ideas, or behaviors when it is asked to generate inappropriate content (Wu et al., 2023c). One potential simple defense strategy is to identify the presence of “red-flagged” keywords (Wei et al., 2023) which strictly violates the usage policies of the LLM vendors, e.g., OpenAI (Deng et al., 2023). Recently, Zhang et al. proposed a goal prioritization technique that prioritizes generating harmless responses over helpful ones at the inference phase (Zhang et al., 2023). This technique can significantly reduce the success rate of jailbreaking attacks in ChatGPT and Llama-2. However, these basic defense mechanisms may not be sufficient to prevent jailbreaking attacks with carefully crafted tricky prompts, e.g., privilege escalation. Furthermore, preventing these attacks still poses a significant challenge because the effective defenses may impair model utility (Li et al., 2023d). By far, SmoothLLM is an effective defense strategy against existing jailbreaking attacks proposed by Zou et al. (Zou et al., 2023). It can be applied for mitigating the instruction-based attacks, e.g., AutoDAN (Liu et al., 2023c), DAN (Shen et al., 2023), and PAIR (Chao et al., 2023). The fundamental concept of SmoothLLM is partially inspired by the randomized smoothing within the adversarial robustness community (Cohen et al., 2019). It involves a two-step process. First, copies of a specified input prompt are duplicated and perturbed. Subsequently, the outputs produced for each perturbed copy are aggregated. The desiderata encompass four key properties: attack mitigation (reduces attack success rate 100 times and 50 times for Llama-2 and Vicuna respectively), non-conservatism, efficiency (in terms of computational resources), and compatibility (different LLMs). For preventing the advanced attack methods, such as multi-step jailbreaking (MSJ) (Li et al., 2023b), and MasterKey (Deng et al., 2023), unsafe prompt detection and filtering technique may effectively refuse to provide the inappropriate response intended by the adversary. On the other hand, self-reminder-based methods (Wu et al., 2024) might create robust defense against more complex attacks such as DeepInception (Li et al., 2023f). However, the efficacy of these mitigation techniques has not yet been evaluated based on these specific attack methods mentioned. Meta introduced Llama Guard, an input-output tool to safeguard LLMs (Inan et al., 2023), designed to effectively mitigate jailbreaking attacks. Llama Guard incorporates a safety risk taxonomy and the applicable policy for data collection and training the tool. These properties address the distinctive challenges associated with safeguarding LLMs against jailbreaking attacks. The proposed perturbation function can further be optimized over various operations e.g., insertion and swaps to make stronger defenses. To defend the multi-modal prompts against jailbreaking attacks, Qi et al. recently proposed DiffPure (Nie et al., 2022), a diffusion model-based countermeasure against the visual jailbreaking examples.
@@
Defense Against Backdoor Attack. Most of the existing research, such as the removal of backdoors by fine-tuning (Sha et al., 2022), model pruning (Liu et al., 2018a), and detecting backdoors by inspecting activations (Chen et al., 2018) are based on backdoor defenses in the white-box setting. Fine-mixing is a mitigation approach designed to prevent backdoors in fine-tuned LMs. It utilizes pre-trained weights through two complementary techniques: (i) a two-step fine-tuning procedure that first combines backdoored weights that have been optimized using pre-trained weights on poisoned data, and then refines the combined weights on a small collection of clean data; (ii) an Embedding Purification (E-PUR) method, addressing potential backdoors in word embeddings (Zhang et al., 2022a). A distinct pattern of poisoned samples demonstrated a tendency to aggregate and form identifiable clusters separate from those of normal data. Building upon this observation, a defense technique named CUBE has been proposed (Cui et al., 2022). It utilized a density clustering algorithm called HDBSCAN to accurately discern clusters within datasets and distinguish the poisoned samples from clean data. By leveraging the capabilities of HDBSCAN (McInnes et al., 2017), CUBE aims to provide an effective means of differentiating clusters associated with both normal and poisoned data. Strategies to defend the backdoor attacks in the black-box setting are still lacking (Kandpal et al., 2023). Perturbation-based and perplexity-based defense methods are also adopted in the literature (Mozes et al., 2023; Yang et al., 2021b; Qi et al., 2020), e.g., RAP which leverages word-based Robustness-Aware Perturbation (RAP) to identify poisoned samples (Yang et al., 2021b) and ONION which eliminates trigger words via empirical analysis of sentence perplexities (Qi et al., 2020). These two methods can be applied to defend against manually designed backdoor triggers such as ProAttack (Zhao et al., 2023a). Masking-Differential Prompting (MDP) serves as an efficient, lightweight, and adaptable defense method against backdoor attacks in prompt-based language models (PLMs), particularly in few-shot learning settings (Xi et al., 2023b). MDP exploits the observation that poisoned samples exhibit increased sensitivity to random masking compared to clean samples. When the trigger is (partially) masked, the language modeling probability of a poisoned sample tends to exhibit significant variations. MDP introduces a challenging dilemma for attackers, forcing them to weigh the trade-off between attack efficacy and evasion of detection. However, MDP falls short on several other PLMs (e.g., GPT-3 (Brown et al., 2020)) and NLP tasks (e.g., paraphrases and sentence similarity (Gao et al., 2020)). Moreover, further studies can be performed to evaluate the performance of random masking-based defense in PLMs when the available data are even scarcer (e.g., one or zero-shot settings (Vinyals et al., 2016), (Wang et al., 2019b)). Again, MDP was proven to be effective for the earlier backdoor attacks, however, it might not safeguard some advanced attacks such as BadPrompt (Cai et al., 2022) and BToP (Xu et al., 2022). Leveraging techniques such as knowledge distillation (Hinton, 2015), outlier-filtering (Xu et al., 2022), and fine-pruning (Liu et al., 2018a) requires further exploration to mitigate the impacts of backdoor attacks, such as BadPrompt (Cai et al., 2022), and BTop (Xu et al., 2022). In order to defend against clean label backdoor triggers, such as LLMBkd (You et al., 2023), a trigger classification approach called REACT (You et al., 2023) has performed significantly better than ONION (Qi et al., 2020) and RAP (Yang et al., 2021b). For the reasoning-based backdoor attacks, e.g., BadChain (Xiang et al., 2024a), shuffling model input might be a potential defense mechanism. Additionally, backdoors detection (Xiang et al., 2024b) (CBD), including backdoor patterns in training data to ensure robustness (Weber et al., 2023) (RAB) can be used to prevent backdoor attacks. Detecting anomalies in input data and parameter decontamination are also used to prevent the attacks executed during the training or fine-tuning phase, e.g., BadAgent (Wang et al., 2024a). Though these defense techniques are primarily proposed for detecting backdoors and mitigating the effects of attacks, most of them are evaluated either on LMs or in other domains, e.g., computer vision. The impact of these mitigation techniques has not yet been evaluated on recently evolved LLMs, e.g., GPT-4 and Llama-3.
@@
Defense Against Data Poisoning Attack. Few solutions exist to defend against poisoning attacks in LLMs. In general, techniques such as data validation, filtering, cleaning, and anomaly detection have been used to protect ML models from poisoning attacks (Pathak, 2023). A detection and filtering approach was designed to identify and filter poisoned data which is collected for performing supervised learning (Baracaldo et al., 2017). Empirical assessments have demonstrated that limiting the number of training epochs is a straightforward method for LMs to reduce the impact of data poisoning such as RoBERTa (Wallace et al., 2020). Identifying poison examples using perplexity can be another technique for small GPT-2 model (Radford et al., 2019) for sentiment analysis tasks. However, it may not identify poisons effectively, specifically, after inspecting the training data, less than half of the poisoned examples can be identified. Identifying these poisoned examples by BERT embedding distance is another method for defending this attack (Wallace et al., 2020). Filtering poisoned samples during training is also used in LMs to defend against data poisoning attacks (Wan et al., 2023). Yan et al. proposed a framework called ParaFuzz for poisoned sample detection for NLP models that leverages the interpretability of model predictions (Yan et al., 2024b). It adopts a software testing technique called fuzzing to distinguish poisoned samples from clean samples. The poisoned data points are often outliers in the training data distribution. Compared to benign training data, a model requires more time to learn the features of poisoned data. To defend against trojaning attacks in LMs, like, TROJANLM (Zhang et al., 2021b), one possible defense technique is to adopt existing techniques from other domains such as images, e.g., STRIP (Gao et al., 2019), detecting trigger-embedded inputs at inference phase (Chen et al., 2018), (Chou et al., 2018), and finding suspicious LMs and retrieving triggers during the model evaluation phase (Wang et al., 2019a), (Chen et al., 2019). Dataset curation techniques, such as removing near-duplicate poisoned samples, known triggers and payloads, and identifying anomalies, can help defend against the attacks performed by manually inserted poisoned samples in training data (Continella et al., 2017). These methods are potentially effective against poisoning attacks, e.g., you auto-complete me (Schuster et al., 2021), AutoPoison (Shu et al., 2023), and TrojanPuzzle attack (Aghakhani et al., 2023). Fine-pruning can also be used as a defense against these attacks (Liu et al., 2018a). For the white-box attacks such as AgentPoison (Chen et al., 2024b), perplexity filtering and query rephrasing (Kumar et al., 2023) are utilized in the literature. On the other hand, advanced attacks, such as NightShade (Bowman, 2023), need advanced defense methods.
@@
The aforementioned defense strategies are mostly designed for LMs, but some of them can be potentially applied to LLMs as well, however, it still lacks in-depth research studies to develop efficient defense techniques to protect LLMs from data poisoning attacks. Moreover, empirical reports have shown that LLMs are becoming more vulnerable to data poisoning attacks, where defenses based on filtering data or lowering model capacity only offer minimal protection at the cost of reduced test accuracy (Wan et al., 2023). Therefore, it requires effective defense methods that can make trade-offs between model utility and the capability of protecting LLMs from data poisoning attacks.